Use deductive reasoning to show that the following procedure produces a number that is three times

the original number. Procedure: Pick a number. Multiply the number by 6, add 10 to the product, divide

the sum by 2, and subtract 5. Hint: Let n represent the original number.


1
Expert's answer
2022-05-11T09:25:47-0400

Let n represent the original number.

Multiply the number by 6. It is 6n

Add 10 to the product. It is 6n+10

Divide the sum by 2. It is (6n+10)/2=3n+5

Subtract 5. It is 3n+5-5=3n

3n/n=3, where 3n is three times

the original number.